OpenStudy (anonymous):

Consider the function f(x)=6x^3−45x^4,x>0. Let F(x) be the antiderivative of f(x) with F(1)=0. Then F(x)=

OpenStudy (anonymous):

sorry! its 6/x^3....

OpenStudy (hunus):

\[F'(x)=f(x)\] \[ f(x)=6x^3−45x^4\] \[F(x)=3x^4/2-9x^5+C\] \[F(1)=3(1)^4/2-9(1)^5+C=0\] \[C=15/2\] \[F(x)=3x^4/2-9x^5+15/2\]

OpenStudy (hunus):

lol